### v1

# Towards Trustworthy Dermatology MLLMs: A Benchmark and Multimodal Evaluator for Diagnostic Narratives

Abstract
Multimodal large language models (LLMs) are increasingly used to generate dermatology diagnostic narratives directly from images. However, reliable evaluation remains the primary bottleneck for responsible clinical deployment. We introduce a novel evaluation framework that combines DermBench, a meticulously curated benchmark, with DermEval, a robust automatic evaluator, to enable clinically meaningful, reproducible, and scalable assessment. We build DermBench, which pairs 4000 real-world dermatology images with expert-certified diagnostic narratives and uses an LLM-based judge to score candidate narratives across clinically grounded dimensions, enabling consistent and comprehensive evaluation of multimodal models. For individual case assessment, we train DermEval, a reference-free multimodal evaluator. Given an image and a generated narrative, DermEval produces a structured critique along with an overall score and per-dimension ratings. This capability enables fine-grained, per-case analysis, which is critical for identifying model limitations and biases. Experiments on a diverse dataset of 4500 cases demonstrate that DermBench and DermEval achieve close alignment with expert ratings, with mean deviations of 0.251 and 0.117 (out of 5) respectively, providing reliable measurement of diagnostic ability and trustworthiness across different multimodal LLMs.
1 Introduction
Dermatologic diseases are prevalent across populations and impose a persistent clinical burden (Karimkhani et al., 2017; Urban et al., 2021). The widespread availability of smartphone and clinical imaging has created a rich visual substrate for computational analysis (Bourkas et al., 2023; Farr et al., 2021; Ouellette & Rao, 2022). Recent advances in multimodal LLMs (Zhou et al., 2024; Nasir, 2025) have made it possible to transform a single dermatology image into a full diagnostic narrative (Yan et al., 2025; Zeng et al., 2025; Lin et al., 2025) that describes visible findings, articulates reasoning, and proposes differential diagnoses (Sellergren et al., 2025; Liu et al., 2023; Zhu et al., 2023; Hurst et al., 2024; Li et al., 2023a; Singhal et al., 2025). This capability is attractive for scalable decision support (Yang et al., 2025; Han et al., 2023; Güneş et al., 2025; Gabashvili, 2023; Pillai et al., 2024; Liu et al., 2024), yet it is also fragile (Asgari et al., 2025; Nakaura et al., 2025; Zack et al., 2024). A narrative that is fluent but clinically unsound can mislead users and create safety risks (Asgari et al., 2025; Chustecki, 2024). Responsible deployment therefore requires rigorous and clinically grounded evaluation of generated texts (Aljamaan et al., 2024).
Evaluating diagnostic narratives is more demanding than producing them (Asgari et al., 2025; Yu et al., 2023a; Liang et al., 2022). Generic text similarity metrics and open ended preference judgments correlate weakly with clinical utility because they ignore visual evidence, downplay patient safety, and fail to test the internal consistency between observed signs and proposed conclusions (Zhang et al., 2019; Delbrouck et al., 2022; Dawidowicz et al., 2024; Trienes et al., 2023; Artsi et al., 2025). Existing resources in dermatology focus on images with categorical labels, while standardized diagnostic narratives and multimodal benchmarks remain scarce. Expert grading is reliable, yet the associated cost limits scale and hinders timely model iteration. These gaps motivate a principled evaluation space that reflects clinical practice and a mechanism that can compare models fairly and monitor risk over time (Wang et al., 2021; Zhang et al., 2023).
We first construct a dataset intended to provide certified exemplars and supervision signals for evaluation. We construct paired image and reasoning data through a dual stream process and obtain clinician curated reference narratives together with concise rationales for scoring. The procedure yields high-quality exemplars and a diverse set of non perfect cases that reveal typical failure modes and provide supervision signals for evaluation.
Figure 1: Overview of DermBench and DermEval. DermBench evaluates a candidate diagnostic narrative by comparing it to a physician-approved reference text for the same image, whereas DermEval evaluates directly from the given image and diagnostic text without requiring a reference.
We then establish a systematic evaluation framework that comprises DermBench and DermEval, as shown in Figure 1. DermBench pairs fixed images with physician approved exemplars and uses an LLM-based judge to compare candidate narratives against certified references. It then assigns scores along clinically grounded dimensions, enabling controlled evaluation of the diagnostic generation capabilities of multimodal models. DermEval is a multimodal evaluator that maps an image and a narrative to scalar scores and a structured critique. At inference time, the model requires no reference text. Given only an image and a diagnostic narrative, it produces an evaluation narrative and per case scores that enable routine assessment. Training proceeds in two stages. In the first stage, the model learns a canonical, machine‑parsable format via token‑level cross‑entropy. In the second stage, it aligns predicted scores with physician ratings using a reinforcement objective with an exponential moving‑average baseline, restricting policy gradients to generated tokens. We adopt six dimensions that capture essential requirements, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision.
Our key contributions are:
  • This study establishes the first benchmark for trustworthy dermatology diagnostic narratives and develops the first dedicated multimodal evaluator for dermatologic diagnosis, enabling fine-grained comparison of model performance across clinical dimensions with explicit attention to trustworthiness.
  • We introduce DermBench, a dermatology benchmark for image-to–diagnostic-narrative generation that includes clinician-certified reference narratives and a fixed six-dimension judging protocol, enabling consistent and transparent comparison across models under the same images and standard exemplars.
  • We develop DermEval, a multimodal evaluator that generates structured rationales and numeric scores and that is trained to align with physician judgments, supporting scalable assessment without requiring reference texts at inference and allowing case-level evaluation.
  • A systematic empirical study and error analysis that reveal model differences across clinical dimensions with an explicit focus on fairness and safety.
2 Related Work
2.1 General and Domain-Specific Multimodal Models
General-purpose VLMs demonstrate strong multimodal understanding across tasks including multi-image reasoning and long-context comprehension (Bai et al., 2025; Li et al., 2024). Parallel advances in “slow-thinking” paradigms highlight explicit multi-step inference: OpenAI o1 employs RL-based deliberation training (Jaech et al., 2024), DeepSeek-R1 scales RL to induce reasoning behaviors without heavy SFT (Guo et al., 2025), and Vision-R1 extends this to multimodality with a curated CoT dataset and RL fine-tuning(Huang et al., 2025).
In the medical domain, LLaVA-Med adapts biomedical figure–caption corpora (Li et al., 2023b). HuatuoGPT-Vision scales medical visual alignment with improvements on health benchmarks (Chen et al., 2024). Med-PaLM M explores generalist multimodal modeling across clinical modalities (Tu et al., 2024). Domain-specific dermatology foundation models further push scale: PanDerm achieves state-of-the-art performance on 28 benchmarks including clinician reader studies (Yan et al., 2025), while DermINO leverages 432K hybrid-pretrained skin images for tasks such as classification, captioning, and fairness evaluation (Xu et al., 2025). Yet across both general-purpose and domain-specific models, the absence of dermatology CoT supervision limits their ability to produce reliable, structured reasoning for clinical workflows.
2.2 Evaluators and Model-as-Judge Approaches
Model as judge has become a central paradigm for evaluating long-form multimodal reasoning because it scales and supports open ended outputs under rubric driven prompts. Benchmarks such as (Yu et al., 2023b) and (Yue et al., 2024) adopt model graded or assisted pipelines. LLaVA-Critic formalizes a general purpose evaluator for open ended vision language outputs (Xiong et al., 2025). In the text setting, (Zheng et al., 2023) analyzes agreement with humans and documents systematic biases, while (Kim et al., 2023) trains an open evaluator that follows detailed rubrics.
Despite this progress, reliability concerns persist. (Gu et al., 2024) highlights instability, sensitivity to prompt length, and misalignment with domain experts. (Shi et al., 2024) quantifies ordering effects in pairwise judgments, and (Szymanski et al., 2025) discusses threats to validity across benchmarking settings. In high stakes medical domains, task specific evaluators grounded in human ratings are therefore recommended. Our work follows this direction by training a dermatology specific evaluator aligned to physician scores and by pairing it with a benchmark that uses clinician certified references.
Figure 2: Dataset construction. A dual stream pipeline is used. The first stream produces clinician verified, high-quality diagnostic narratives that become the certified references for DermBench. The second stream produces diagnostic narratives of varying quality. Image and text pairs from both streams are used to train the evaluator DermEval.
3 Methodology
In this section we present an integrated pipeline that spans data construction, benchmarking, and evaluator learning. We first describe the dataset making process in Section 3.1, which builds paired image and reasoning data through a dual stream procedure that uses structured prompts for caption generation and hierarchical clinical reasoning, followed by expert curation that produces certified references and evaluation rationales for each image and diagnostic narrative. We then define DermBench in Section 3.2, which pairs each image with its certified reference and employs an LLM judge to score candidate narratives on six dimensions, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, thereby establishing the official metrics for reporting. Finally we introduce DermEval in Section 3.3, a LLaVA based evaluator that maps an image and a narrative to six scalar scores and a structured evaluation. Training proceeds in two stages. Stage one learns a canonical and parsable evaluation format with token level cross entropy. Stage two aligns emitted scores to physician scores using score oriented REINFORCE with an exponential moving average baseline while sending policy gradients only to generated tokens. Section 3.4 specifies the unified objective that combines the text loss and the reinforcement loss with fixed weights and summarizes the inference time prompting and score extraction protocol.
3.1 Dataset Construction
We construct the dataset through two coordinated streams, as shown in Figure 2. In the high-quality stream we first select a large set of dermatology images with diverse and balanced categories. For each image we use Gemini 2.5 Pro (Comanici et al., 2025)to prompt a vision language model to produce a caption that reports only observable findings such as anatomic location, lesion morphology, and surface characteristics while forbidding diagnostic speculation. The captioning prompt is: “Please briefly describe the physical location of the affected skin and the observable medical features of the skin condition. Do not make any differential diagnosis. Start your response with ‘This image shows …’.” A second LLM then receives the caption together with the ground truth label and is instructed to simulate clinical reasoning. The hierarchical reasoning prompt is: “Simulate expert hierarchical reasoning for dermatologic diagnosis using the provided caption and the ground truth {DISEASE_NAME}. Begin with high level categorization, progressively refine to specific diseases and pathological features, and conclude with a coherent diagnostic judgment.” The model reasons stepwise and presents the disease name only in the final sentence. We normalize the output into a hierarchical chain of thought that states coarse disease families, then intermediate descriptors, and finally the specific diagnosis. Board certified dermatologists rate every text on six axes from 0 to 5, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. If any dimension receives a score below 5, clinicians revise the text until all dimensions are rated 5, after which the recorded scores are fixed at 5.
The regular stream reuses the same images. For each image we again obtain a caption without diagnostic content using Gemini 2.5 Pro and we reuse the same captioning prompt. A second LLM receives the caption and the image without the label and is instructed to diagnose through stepwise reasoning, placing the inferred disease name in the final sentence. We reuse the same hierarchical reasoning prompt for this stream. The output is converted to the same hierarchical chain of thought format. Board certified dermatologists score each text using the same six criteria on a scale from 0 to 5.
High score texts constitute the benchmark introduced in Section 3.2. For each image, its diagnostic narrative, and the associated six‑dimension scores, we additionally elicit an evaluation rationale using the following instruction: “Given the dermatology image, the generated diagnostic narrative, and the six numeric scores for Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision, produce a concise justification for each dimension. Structure the output as six titled sections that match the dimension names. In each section restate the score, cite concrete evidence from the narrative or observable findings that supports the score, and end with one actionable suggestion for improvement. Do not propose a new diagnosis and do not alter the scores.” The resulting rationale is saved as the evaluation text label. All image–diagnostic text–evaluation text triples, including both high and low score cases, are used to train DermEval in Section 3.3.
Figure 3: DermBench evaluation workflow. A candidate LLM generates a diagnostic narrative from a standardized prompt. An LLM judge compares the narrative with the clinician-certified reference and assigns six scores, namely Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision.
Figure 4: DermEval training pipeline. The evaluator takes an image and a diagnostic text, generates a structured evaluation, and an external LLM extracts six scores in the range from zero to five. Physician scores define a negative mean squared error reward, an exponential moving average baseline yields a low variance advantage, and policy gradients are applied only to the generated segment.
3.2 DermBench Construction and LLM Scoring
We construct DermBench from the images and certified high-quality reference texts defined in the previous subsection, as shown in Figure 3. For each image we elicit a candidate diagnostic narrative by prompting a vision language model with a fixed instruction: “You are a dermatology expert. Analyze the image by describing key features such as morphology, color, distribution, and surface changes. Interpret the findings with clinical reasoning. Provide two to three differential diagnoses. Justify the top diagnosis. Conclude with ‘The answer is {DISEASE_NAME}’ where {DISEASE_NAME} is the disease name.” The resulting narrative is treated as the candidate text. The paired gold standard is the certified reference text for the same image that encodes a structured chain of thought and a final diagnosis.
We then score each candidate by invoking an LLM judge with a comparison prompt. The judge receives the candidate text and the gold reference text and is instructed to assign a score from 0 to 5 on six dimensions, specifically Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. The comparison instruction states: “Given the two passages above, where the first is our generated diagnostic text and the second is the gold standard reference, compare them and assign our generated text a score from 0 to 5 for Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Use 0 for the lowest and 5 for the highest.” The six scores constitute the official DermBench metrics.
3.3 DermEval Training with Score Oriented REINFORCE with EMA Baseline
We finetune a LLaVA model into DermEval that maps an image and a diagnostic text to six scalar scores and a structured evaluation text, as shown in Fig 4. The six dimensions are Accuracy, Safety, Medical Groundedness, Clinical Coverage, Reasoning Coherence, and Description Precision. Physician scores on these six dimensions are normalized to the interval from zero to five. Training proceeds in two stages. Stage 1 teaches the model to produce a canonical and parsable evaluation format. Stage 2 aligns the scores emitted in the generated evaluation text with physician scores through reinforcement learning.
In Stage 1 the model receives an image $I$ and a diagnostic text $d$. The target output is a templated evaluation text $y^{*}$ that contains six explicit score fields. The model is optimized with token level cross entropy to obtain stable structured outputs that can be reliably parsed in the next stage
$\mathcal{L}_{\mathrm{TEXT}}=\mathrm{CE}\!\left(y,\,y^{*}\right).$ (1)
where $x$ denotes the prompted input that is composed by the conversation template and the image sentinel with $(I,d)$. The variable $y$ is the predicted evaluation text and $y^{*}$ is the target template.
In Stage 2 we introduce Score Oriented REINFORCE with exponential moving average (EMA) Baseline. For each instance we construct the same prompt as in inference and generate a single evaluation text $\hat{y}$ without gradients. An external LLM parses $\hat{y}$ to extract the six scores $\hat{\mathbf{s}}\in[0,5]^{6}$, with the prompt for the external LLM parser shown in Figure 4. Let $\mathbf{s}^{*}\in[0,5]^{6}$ be the physician scores, and let $\mathcal{I}_{y}$ be the set of indices that are successfully parsed with size $K_{y}$. The numeric alignment reward is the negative mean squared error on valid dimensions
$r(\hat{\mathbf{s}},\mathbf{s}^{*})\;=\;-\,\frac{1}{K_{y}}\sum_{k\in\mathcal{I}_{y}}\big(\hat{s}_{k}-s_{k}^{*}\big)^{2}.$ (2)
where $\hat{\mathbf{s}}$ denotes the parsed scores, $\mathbf{s}^{*}$ denotes the physician scores, $\mathcal{I}_{y}$ is the valid index set, and $K_{y}=|\mathcal{I}_{y}|$. If a metric score cannot be parsed, it is excluded from the loss.
We maintain an exponential moving average baseline $b$ with momentum $\beta\in(0,1)$ in order to reduce the variance of the REINFORCE gradient and to stabilize updates across iterations. The baseline aggregates recent rewards with a decaying weight so that it tracks the reward trend while remaining robust to outliers. This yields a low noise estimate of the expected reward without requiring an additional greedy generation. The advantage is then computed as the deviation of the current reward from this baseline
$b\leftarrow\beta\,b+(1-\beta)\,r,\quad\mathrm{adv}\;=\;r-b.$ (3)
where $b$ is the moving baseline, $\beta$ is the momentum coefficient, $r$ is the reward, and $\mathrm{adv}$ is the advantage.
To send gradients only to generated tokens we run a teacher forcing pass on the sampled sequence and accumulate the log likelihood over the generation segment $\mathcal{T}$ that starts after the prompt. The reinforcement objective is
$\mathcal{L}_{\mathrm{RL}}\;=\;-\,\mathrm{adv}\cdot\operatorname{mean}_{t\in\mathcal{T}}\log p_{\theta}\!\left(y_{t}\mid y_{<t},x\right).$ (4)
where $\mathcal{T}$ denotes the set of generated token positions and $p_{\theta}$ denotes the model distribution.
Stage 2 minimizes
$\mathcal{L}\;=\;\lambda_{\mathrm{RL}}\,\mathcal{L}_{\mathrm{RL}}\;+\;\lambda_{\mathrm{TEXT}}\,\mathcal{L}_{\mathrm{TEXT}},$ (5)
where $\lambda_{\mathrm{RL}}>0$ and $\lambda_{\mathrm{TEXT}}>0$ are the weights for the reinforcement loss and the text loss, and $\mathcal{L}_{\mathrm{TEXT}}$ and $\mathcal{L}_{\mathrm{RL}}$ are defined above.
3.4 Implementation Details
For building DermEval, training is conducted on eight NVIDIA RTX 4090 GPUs. The backbone is LLaVA v1.5 with 7B parameters. Finetuning uses LoRA with rank set to 64, alpha set to 16, dropout set to 0.05, an empty LoRA weight path, and LoRA bias set to none. Both stages use the same optimization schedule. The warmup ratio is 0.03. The initial learning rate is 2e-5 and the scheduler follows a cosine decay. Each stage is trained for five epochs. Training uses 2,000 certified exemplar images with perfect scores and 2,000 imperfect low-score images. In Stage 2 the loss uses a fixed weighting where the reinforcement learning loss has weight 0.5 and the text loss has weight 1.0. The exponential moving average baseline in Stage 2 uses a beta of 0.9.
4 Experiments
4.1 Alignment Tests with Expert Annotations
To verify the alignment of DermBench and DermEval with expert judgment, we selected nine representative models spanning three categories, namely general purpose multimodal models including Gemini-2.5-Flash, GPT-4o mini, and GPT-o4 mini, reasoning-focused multimodal models including Vision-R1, InternVL3, LLaVA-o1, and Qwen2.5-VL, and medical domain models including HuatuoGPT-Vision and LLaVA-Med-7B. We evaluated 500 images sampled from DermNet that were not used for DermEval training to avoid overfitting. For each image and for each model we generated a diagnostic narrative under the standardized instruction protocol defined in Section 3, yielding 4500 narratives in total. Dermatologists then assigned scores on six clinically grounded dimensions for every narrative. In addition, for each image we produced a perfect reference narrative following the procedure in Section 3.1. We subsequently evaluated all model generated narratives with DermBench using the certified reference and with DermEval without any reference. Finally we computed the per dimension errors of DermBench and DermEval relative to physician scores. All scores are rounded to the nearest integer on the 0 to 5 scale in order to align with physician ratings. Results are reported in Table 1.
Table 1: Mean absolute error (MAE) between model scores and expert annotations across six metrics. Metric abbreviations: Acc (Accuracy), Safe (Safety), MedG (Medical Groundedness), Cover (Clinical Coverage), Reason (Reasoning Coherence), Desc (Description Precision).
Evaluator Acc Safe MedG Cover Reason Desc
DermBench 0.251 0.314 0.369 0.456 0.412 0.377
DermEval 0.117 0.230 0.176 0.152 0.236 0.147
Both DermBench and DermEval track physician judgments closely, with all errors below 0.5 on the 0 to 5 scale. The macro average MAE is 0.36 for DermBench and 0.18 for DermEval, which indicates low discrepancy overall and a consistent advantage for DermEval. DermEval achieves lower MAE on every dimension, with the largest reductions on Clinical Coverage and Description Precision, while DermBench remains competitive as a reference-anchored comparator. These findings support DermBench for controlled comparisons against certified references and DermEval for case level scoring without references, and show that the two evaluators provide reliable and complementary mechanisms for measuring diagnostic quality. Additionally, we present examples in the Supplementary Material showing that the scores from DermBench and DermEval closely match physician ratings.
4.2 Evaluating Multimodal Models
Having established the alignment of DermBench and DermEval with expert judgments, we benchmark nine multimodal models on the DermNet dataset, namely Gemini-2.5-flash, GPT-4o-mini, GPT-o4mini, Vision-R1, InternVL3, LLaVA-o1, Qwen2.5-VL, HuatuoGPT-Vision, and LLaVA-Med-7B. For every image in DermNet each model generates a diagnostic narrative under the standardized instruction protocol defined in Section 3. To supply certified references for controlled comparisons we additionally select 4000 images from DermNet and create perfect reference narratives following the procedure in Section 3.1. The distribution of disease categories used in the benchmark is shown in Figure 5. We then evaluate the narratives produced by all nine models using both DermBench and DermEval. For each method and each model we compute scores on the six clinically grounded dimensions and report the mean values. Results are shown in Table 2.
Figure 5: Distribution of skin disease categories covered by our dataset. The pie chart illustrates the proportion of images in each of the 23 dermatological categories used for model training and evaluation.
Table 2: Benchmarking 9 LLMs on the DermBench and DermEval evaluators. Each model is evaluated across six standard clinical metrics under each setting.
Model DermBench DermEval
Acc Safe MedG Cover Reason Desc Acc Safe MedG Cover Reason Desc
Gemini-2.5-Flash(Comanici et al., 2025) 3.143 4.089 3.439 3.974 3.990 4.339 3.201 4.128 3.482 3.952 4.031 4.298
InternVL3(Zhu et al., 2025) 2.408 3.288 2.505 3.112 3.121 3.750 2.369 3.251 2.563 3.074 3.163 3.706
Qwen2.5-VL(Bai et al., 2025) 1.787 2.658 1.914 2.609 2.565 3.328 1.824 2.701 1.872 2.556 2.617 3.371
GPT-4o-mini(Hurst et al., 2024) 3.134 4.085 3.520 3.987 4.023 4.483 3.182 4.042 3.566 4.021 3.993 4.451
Vision-R1(Huang et al., 2025) 2.337 3.353 2.492 2.868 2.798 3.340 2.301 3.312 2.541 2.826 2.839 3.287
LLaVA-o1 (Xu et al., 2024) 2.667 3.807 2.828 2.840 3.237 3.750 2.619 3.764 2.881 2.793 3.198 3.721
GPT-o4-mini(Jaech et al., 2024) 2.602 4.291 3.564 3.300 3.697 4.079 2.659 4.249 3.523 3.347 3.741 4.126
HuatuoGPT-Vision(Chen et al., 2024) 2.483 3.406 2.592 3.323 3.330 3.942 2.436 3.462 2.548 3.367 3.291 3.983
LLaVA-Med-7B(Li et al., 2023b) 1.207 2.199 1.220 1.388 1.452 2.531 1.169 2.157 1.269 1.431 1.496 2.487
Figure 6: Average performance of nine multimodal models across six clinical metrics, computed by averaging DermBench and DermEval scores.
DermBench and DermEval yield highly consistent scores across all six metrics. For the same model and metric, the two evaluators typically differ by no more than 0.05, which indicates that the reference-free DermEval closely tracks the reference anchored DermBench. The relative ordering of systems is preserved. For example, Gemini-2.5-flash and GPT-4o-mini are the top performers on Accuracy under both evaluators with values around 3.1 to 3.2, while LLaVA-Med-7B remains the lowest near 1.2. On Safety, GPT-o4-mini ranks first under both settings with scores around 4.25, and the same models form the upper tier for Medical Groundedness and Coverage with only minor swaps in the top position. Reasoning and Description Precision show the same pattern, with Gemini-2.5-flash and GPT-4o-mini alternating at the top and LLaVA-Med-7B consistently at the bottom. These concordant trends demonstrate stable model ordering and support the conclusion that DermEval has learned physician-aligned scoring behavior, making it a reliable judge for case-level assessment.
5 Discussion
This work introduces a clinically grounded evaluation infrastructure for image to diagnostic narrative generation in dermatology. DermBench enables controlled comparisons by pairing fixed images with physician approved references and by using an LLM-based judge to assign scores on six dimensions. DermEval enables reference-free case level assessment by generating structured critiques and scores directly from an image and a narrative. Alignment tests show close agreement with expert annotations, and large scale benchmarking across nine models demonstrates consistent evaluator behavior and a granular view of model strengths and weaknesses.
Our study has several limitations. First, reliance on DermNet may limit real-world diversity; we plan to expand to multi-source images and perform cross-site validation. Second, certified references are model-drafted and clinician-revised, which can introduce bias; to mitigate this we will diversify templates and conduct blinded multi-rater editing. Third, physician scoring may be noisy; we will stabilize ratings with anchor items, double scoring with adjudication, and uncertainty reporting for difficult cases.
6 Conclusion
In this paper, we presented DermBench and DermEval as a clinically grounded evaluation infrastructure for image to diagnostic narrative generation in dermatology. DermBench uses physician approved references and an LLM-based judge to score candidate narratives on six dimensions under controlled prompts. DermEval provides reference-free, case-level assessment by producing structured critiques and scores from an image and a narrative. Alignment tests show close agreement with expert annotations across all metrics. Benchmarking nine contemporary multimodal models on DermNet reveals consistent evaluator behavior and exposes complementary strengths and weaknesses by dimension. Future work will extend the benchmark and the evaluator to larger and more diverse corpora, refine error taxonomies and calibration, and study integration with prospective human oversight in clinical workflows.